OpenStudy (anonymous):

use the sum/difference identities to evaluate exactly cos 105 degree Help plzz :(

OpenStudy (anonymous):

105 = 60 + 45

OpenStudy (phi):

105= 60+45 do you know cos(A+B)= ?

OpenStudy (anonymous):

cos(1/2+\[\sqrt{2}/2\])

OpenStudy (phi):

there is a formula for the cosine of the sum of two angles cos(A+B)= cos(A)cos(B) - sin(A)sin(B) in your problem, A=60 and B=45 can you finish?

OpenStudy (anonymous):

\[1/2*\sqrt{2}/2+\sqrt{3}/2*\sqrt{2}/2\]

OpenStudy (phi):

now simplify

OpenStudy (anonymous):

\[1/\sqrt{2}\]

OpenStudy (phi):

but notice the two terms are subtracted (not added)

OpenStudy (anonymous):

[-1/\sqrt{2}\]

OpenStudy (phi):

you should start with \[ \frac{1}{2}\cdot \frac{\sqrt{2}}{2} - \frac{\sqrt{3}}{2} \frac{\sqrt{2}}{2} \]

OpenStudy (anonymous):

\[\sqrt{2}/4-\sqrt{6}/4\]

OpenStudy (phi):

yes, or (which is the same thing) \[ \frac{\sqrt{2}-\sqrt{6}}{4} \] or \[ \frac{\sqrt{2}(1-\sqrt{3})}{4}\] they all evaluate to the same number. You can use a calculator and see cos(150 deg) = -0.258819...

OpenStudy (phi):

* cos(105)= -0.258819...

OpenStudy (anonymous):

thanks

OpenStudy (phi):

yw